# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ 5. Besicovitch Norms of Spectral Sums
We begin by placing the spectral sums in the Besicovitch Hilbert space, which provides the natural inner product for Cesàro-distributed almost periodic functions. The functions $S_{\zeta}$ and $S_{L}$ live in the Besicovitch Hilbert space $B^{2}$ with inner product
@@ 6. Finite Density Bounds via Jacobi–Anger Decomposition
The Jacobi–Anger decomposition makes this obstruction explicit: resonances contribute correction terms to the characteristic function. The key insight of our approach is that these correction terms are self-suppressing: the Bessel factors $J_{n_{k}}(b_{k}t)$ with $n_{k}\neq 0$ decay exponentially in $|n_{k}|$, while the “inactive” $J_{0}$ factors provide polynomial decay in $t$. This suppression is strong enough to guarantee convergence of the density integral regardless of how many resonances exist. The telescoping argument then extends this finite-$M$ bound to the full infinite series by showing that new resonances introduced at each step are controlled by their own Bessel weights.
@@ 6.1. The Jacobi–Anger decomposition
We begin with the decomposition of the Cesàro characteristic function of $S_{L}^{(M)}$.
@@ 6.2. Resonance certification for $M=20$
Let $\gamma_{1}^{\prime},\ldots,\gamma_{20}^{\prime}$ be the first $20$ zero ordinates of $L(s,\chi_{5})$, certified to $70$ decimal places with individual ARB bounds $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{5})|<4\times 10^{-371}$ (Appendices B and E).
@@
    (Large coefficients.) For any integer vector with $\max|n_{k}|\geq 1001$, the Bessel contribution to the density integral is bounded by $10^{-850}$.
Parts (i) and (ii) together cover all nonzero $\mathbf{n}\in\mathbb{Z}^{20}$: part (i) certifies nonexistence for $\max|n_{k}|\leq 1000$ via PSLQ at 70-digit precision, and part (ii) bounds the contribution of any $\mathbf{n}$ with $\max|n_{k}|\geq 1001$ via Bessel decay. Consequently, the density formula
@@
A brute-force search at double precision found no exact resonances among orders-3 and orders-4 combinations and no pairwise relation with $|n_{j}|,|n_{k}|\leq 50$. The nearest misses were $|\gamma_{4}^{\prime}+\gamma_{10}^{\prime}+\gamma_{12}^{\prime}-\gamma_{19}^{\prime}|=0.00037$ (order-4) and $|3\gamma_{3}^{\prime}-\gamma_{14}^{\prime}|=0.0079$ (pairwise), consistent with Proposition 6.3 (i). (The bound $\|\mathbf{m}\|_{1}\geq 184$ mentioned in §A.7 comes from a separate precision-control test at 20-digit input, not the primary 70-digit certification.)
@@ 6.3. Bessel decay estimates
The convergence of the resonance sum relies on the inactive $J_{0}$ factors. For a resonance vector $\mathbf{n}$ with active set $A=\{k:n_{k}\neq 0\}$, the M|A|M-|A| indices not in $A$ contribute $J_{0}(b_{k}t)$ factors to the product. These factors collectively provide exponential suppression.
@@
Supercritical region $2T^{*},\infty)$. For $t\geq 2T^{*}$, each inactive factor satisfies $|J_{0}(b_{k}t)|\leq\sqrt{2/(\pi b_{k}t)}$. Each active factor satisfies the uniform asymptotic $|J_{n_{k}}(b_{k}t)|\leq\sqrt{2/(\pi\sqrt{(b_{k}t)^{2}-n_{k}^{2}})}$, which for $b_{k}t\geq 2|n_{k}|$ is bounded by $\sqrt{2/(\pi b_{k}t)}\cdot(1-(n_{k}/(b_{k}t))^{2})^{-1/4}\leq 2\sqrt{2/(\pi b_{k}t)}$. The total decay is therefore $O(t^{-M/2})$, which is integrable for $M\geq 3$.
@@
Explicit calculation gives
$C_{1}=\frac{T^{*}}{2},\qquad C_{2}=\frac{3T^{*}}{2}\cdot\prod_{k\colon n_{k}\neq 0}\bigl(0.675\,|n_{k}|^{-1/3}\bigr),\qquad C_{3}=\int_{2T^{*}}^{\infty}t^{-M/2}\,dt=\frac{(2T^{*})^{1-M/2}}{M/2-1},$
where $C_{3}$ is finite for $M\geq 3$. Each of $C_{1}$, $C_{2}$, $C_{3}$ depends on $M$ only through the minimal weight $b_{M}=2/(\tfrac{1}{4}+{\gamma_{M}^{\prime}}^{2})$, which decreases monotonically in $M$. Since $b_{M}$ is bounded below by a positive constant for all finite $M$, the three constants are uniformly bounded for $M\geq 3$. ∎
@@
Reproducibility. The values of $I_{\mathbf{n}}$ were computed by ARB quadrature (acb_calc_integrate, 256-bit precision) with the integrand
@@
where $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$ with zeros from Appendix B. For $N\leq 30$ the integration domain [0, 2000][0,\,2000] suffices (the integrand decays as $t^{-M/2}$ for $t\gg N/b_{1}$). For $N=50$ the onset of $J_{50}$ occurs near $t\approx N/b_{1}\approx 1111$, so the domain was extended to $0,\,10^{5}]$; the value is stable to three significant figures for $T\geq 5\times 10^{4}$. The scalings $(e/4)^{N}$ and $N^{-8.67}$ were evaluated directly.
@@ 6.5. Unconditional finiteness at finite $M$
The exponential decay $\eta<1$ comes entirely from the inactive $J_{0}$ factors. Without them, the resonance integral decays only polynomially, which cannot overcome even polynomial lattice counting. The following table illustrates the suppression for a three-active resonance (indices $1,2,3$, each with $|n_{k}|=1$) as $M$ increases:
@@
33 $3.676$ $1.000$
$10$ $1.794$ $0.488$
2020 $1.720$ $0.468$
$30$ $1.706$ $0.464$
Each additional inactive $J_{0}$ factor reduces the resonance integral, converging to a limit as $M\to\infty$. The values of $I_{\mathbf{n}}$ were computed by ARB quadrature (acb_calc_integrate, 256-bit precision, $T=2000$), with three active indices $(n_{1},n_{2},n_{3})=(1,1,-1)$ and the remaining M3M-3 indices set to zero (inactive).
@@ 7.1. Monotonicity of the main term
We begin with the observation that the main-term integral is monotone decreasing in $M$.
@@
The main-term sequence $\{f_{\mathrm{main}}^{(M)}\}$ is therefore a decreasing sequence bounded below by zero, and converges to a finite limit. The values in the following table were computed by rigorous numerical quadrature of $\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$ with $T=2000$ using ARB’s acb_calc_integrate at 256-bit precision. These are the unsigned main-term integrals and serve as upper bounds on $\|f_{S_{L}^{(M)}}\|_{\infty}$; the signed integrals that equal the actual density $f^{(M)}(0)$ under verified resonance absence are smaller (see Remark 7.2).
@@
33 $12.38$ $7.84$
$10$ $10.59$ $8.29$
2020 $10.52$ $8.31$
This is consistent: each additional $J_{0}(b_{k}t)$ factor reduces the absolute value of the integrand but also reduces the negative lobes, pushing the signed integral upward. The finiteness proof uses only the unsigned bound. The monotone increase of the signed integral confirms that $f^{(M)}(0)$ converges to $f_{S_{L}}(0)=8.3129$ from below. Both columns are computed by ARB quadrature at $256$-bit precision with $T=2000$; for any $T\geq 500$ the contribution from $T,\infty)$ is below $10^{-10}$ due to the $t^{-M/2}$ decay of the Bessel product.
@@ 7.2. The self-referential telescoping bound
In the supercritical region, the uniform asymptotic gives $|J_{n_{M+1}}(b_{M+1}t)|\leq\sqrt{2/(\pi\sqrt{(b_{M+1}t)^{2}-n_{M+1}^{2}})}\leq 2\sqrt{2/(\pi b_{M+1}t)}$ for $b_{M+1}t\geq 2|n_{M+1}|$. For the remaining product, we use the uniform envelope bound $|J_{n}(x)|\leq\min(1,\sqrt{2/(\pi|x|)})$, which holds for all orders $n\geq 0$ and all $x>0$. This gives $\prod_{k=1}^{M}|J_{n_{k}}(b_{k}t)|\leq\prod_{k=1}^{M}\min(1,\sqrt{2/(\pi b_{k}t)})=:g_{M}(t)$.
@@
2020 $1.18\times 10^{-1}$ $0.00093$ $128$
$30$ $6.27\times 10^{-2}$ $0.00049$ $127$
The ratio $I_{\mathbf{n}}/b_{M}$ is bounded and stabilizing, confirming $C_{M}\leq C$ uniformly. For $n_{M}=2$, the ratio $I_{\mathbf{n}}/b_{M}^{2}$ similarly stabilizes near $1550$. Each entry $I_{\mathbf{n}}$ in the table corresponds to a resonance vector with $n_{M}=1$ and all other components zero; the integral was evaluated by ARB quadrature (acb_calc_integrate, 256-bit precision, $T=2000$).
@@
Since $\prod|J_{0}|$ is integrable on $0,T^{*}]$ (it is bounded), the true contribution is $O(b_{M+1}^{2})$ with an effective constant given by $\frac{1}{4\pi}\int_{0}^{T^{*}}t^{2}\prod|J_{0}(b_{k}t)|\,dt$, which is finite and bounded uniformly in $M$ (the product can only decrease as $M$ grows).
@@
$10$ $6.70\times 10^{-3}$ $0.00247$ $6.1\times 10^{-6}$ $1100$
$15$ $1.83\times 10^{-3}$ $0.00138$ $1.9\times 10^{-6}$ $966$
2020 $7.91\times 10^{-4}$ $0.00093$ $8.6\times 10^{-7}$ $924$
The ratio $|f^{(M)}(0)-f^{(M-1)}(0)|/b_{M}^{2}$ is bounded and slowly varying, confirming $\alpha=2$. The marginal changes were computed as differences of successive $f^{(M)}(0)$ values obtained by ARB quadrature (acb_calc_integrate, 256-bit precision, $T=2000$); the weights $b_{M}=2/(1/4+{\gamma_{M}^{\prime}}^{2})$ use the verified zero ordinates from the appendix.
@@ 7.3. The major-arc/minor-arc decomposition
The values $\epsilon_{M}=\sum_{k>M}b_{k}$ below are computed from the first 200 certified zeros of $L(s,\chi_{5})$ (Appendix C) plus the analytic tail bound $0.0072$ via Abel summation against the zero-counting formula; see Appendix A for the full methodology.
@@
Under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (in fact $d_{\infty}=0$, verified for $M\leq 20$), the Cesàro distribution of $S_{L}(\delta)=\sum_{k=1}^{\infty}b_{k}\cos(\gamma_{k}^{\prime}\delta)$ has a density at the origin satisfying $f_{S_{L}}(0)<\infty$. The same holds for $f_{S_{\zeta}}(0)$.
@@
It is natural to ask whether the passage from verified resonance absence at finite $M$ to the limiting density $f_{S_{L}}(0)<\infty$ can be achieved by purely approximation-theoretic means, bypassing the hypothesis $d_{\infty}<\infty$. Since $S_{L}^{(M)}\to S_{L}$ in $B^{2}$, the value distributions converge weakly (by Cauchy–Schwarz applied to characteristic functions and Lévy continuity). However, weak convergence does not imply pointwise convergence of densities. The Fourier inversion formula $f^{(M)}(0)=(2\pi)^{-1}\int\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt$ requires the product factorization of the characteristic function, which in turn requires $\mathbb{Q}$-linear independence of the frequencies $\gamma_{1}^{\prime},\ldots,\gamma_{M}^{\prime}$. The dominated convergence argument that yields $f^{(M)}(0)\to f^{(\infty)}(0)$ relies on the monotone bound $|\varphi^{(M+1)}(t)|\leq|\varphi^{(M)}(t)|$, which holds only under the product structure. Without independence, the characteristic function need not factor, the monotone domination fails, and no alternative $L^{1}$ dominator is available. The obstruction is therefore structural, not computational: every Fourier-analytic path to $f_{S_{L}}(0)<\infty$ passes through the independence of zero ordinates.

### 12 Mar 2026 v3

@@ 5. Besicovitch Norms of Spectral Sums
The spectral sums $S_{\zeta}$ and $S_{L}$ live in the Besicovitch Hilbert space $B^{2}$, which provides the inner product for Cesàro-distributed almost periodic functions:
@@ 6. Finite Density Bounds via Jacobi–Anger Decomposition
The Jacobi–Anger decomposition makes this obstruction explicit: resonances contribute correction terms to the characteristic function. These correction terms are self-suppressing: the Bessel factors $J_{n_{k}}(b_{k}t)$ with $n_{k}\neq 0$ decay exponentially in $|n_{k}|$, while the “inactive” $J_{0}$ factors provide polynomial decay in $t$. This suppression is strong enough to guarantee convergence of the density integral regardless of how many resonances exist. The telescoping argument then extends this finite-$M$ bound to the full infinite series by showing that new resonances introduced at each step are controlled by their own Bessel weights.
@@ 6.1. The Jacobi–Anger decomposition
The Cesàro characteristic function of $S_{L}^{(M)}$ admits the following decomposition.
@@ 6.2. Resonance certification for $M=20$
Let $\gamma_{1}^{\prime},\ldots,\gamma_{20}^{\prime}$ be the first $20$ zero ordinates of $L(s,\chi_{5})$, certified to $70$ decimal places with individual ARB bounds $|L(\tfrac{1}{2}+i\gamma_{k}^{\prime},\chi_{5})|<10^{-449}$ (Appendices B and E).
@@
    (Large coefficients.) For any integer vector with $\max|n_{k}|\geq 1001$, the Bessel contribution to the density integral is bounded by $10^{-849.5}$.
Parts (i) and (ii) together cover all nonzero $\mathbf{n}\in\mathbb{Z}^{20}$: part (i) certifies nonexistence for $\max|n_{k}|\leq 1000$ via PSLQ at 70-digit precision, and part (ii) bounds the contribution of any $\mathbf{n}$ with $\max|n_{k}|\geq 1001$ via Bessel decay (ARB-certified total $10^{-849.5}$). Consequently, the density formula
@@
A brute-force search at double precision found no exact resonances among orders-3 and orders-4 combinations and no pairwise relation with $|n_{j}|,|n_{k}|\leq 50$. The nearest misses were $|\gamma_{4}^{\prime}+\gamma_{10}^{\prime}+\gamma_{12}^{\prime}-\gamma_{19}^{\prime}|=0.00037$ (order-4) and $|3\gamma_{3}^{\prime}-\gamma_{14}^{\prime}|=0.0079$ (pairwise), consistent with Proposition 6.3 (i).
@@ 6.3. Bessel decay estimates
The convergence of the resonance sum relies on the inactive $J_{0}$ factors. For a resonance vector $\mathbf{n}$ with active set $A=\{k:n_{k}\neq 0\}$, the M|A|M-|A| indices not in $A$ contribute $J_{0}(b_{k}t)$ factors to the product. These factors contribute polynomial decay in $t$, providing the integrability needed for the resonance sum; the exponential decay in $|n_{*}|$ comes separately from the subcritical bound on the active high-order Bessel factor $J_{n_{*}}(b_{*}t)$.
@@
Supercritical region $2T^{*},\infty)$. For $t\geq 2T^{*}$, each inactive factor satisfies $|J_{0}(b_{k}t)|\leq\sqrt{2/(\pi b_{k}t)}$. Each active factor satisfies the uniform asymptotic
$|J_{n_{k}}(b_{k}t)|\leq\sqrt{\frac{2}{\pi\sqrt{(b_{k}t)^{2}-n_{k}^{2}}}},$
which for $b_{k}t\geq 2|n_{k}|$ is bounded by $\sqrt{2/(\pi b_{k}t)}\cdot(1-(n_{k}/(b_{k}t))^{2})^{-1/4}\leq 2\sqrt{2/(\pi b_{k}t)}$. The total decay is therefore $O(t^{-M/2})$, which is integrable for $M\geq 3$.
@@
Here $C_{1}$, $C_{2}$, $C_{3}$ are positive constants depending on $\{b_{k}\}$, $M$, and $|A|$, each finite for $M\geq 3$. Specifically, $C_{1}$ is of order $T^{*}/2$ after integrating the power-series bound; $C_{2}$ absorbs the interval width $3T^{*}/2$, the Landau factors $0.675\,|n_{k}|^{-1/3}$ for each active index, and the constants $\prod_{k\colon n_{k}=0}(2/(\pi b_{k}))^{1/2}$ from the inactive $\sqrt{2/(\pi b_{k}t)}$ bounds evaluated at the onset $t\sim T^{*}$; and $C_{3}=(2T^{*})^{1-M/2}/(M/2-1)$ after integrating the full $t^{-M/2}$ decay. Since $b_{M}$ is bounded below by a positive constant for all finite $M$, all three constants are uniformly bounded for $M\geq 3$. ∎
@@
Reproducibility. The values of $I_{\mathbf{n}}$ were computed by ARB quadrature via strip decomposition (acb.integral on gap intervals; see §A.6, 256-bit precision) with the integrand
@@
where $b_{k}=2/(\tfrac{1}{4}+{\gamma_{k}^{\prime}}^{2})$ with zeros from Appendix B. For $N\leq 30$ the integration domain [0, 2000][0,\,2000] suffices (the integrand decays as $t^{-M/2}$ for $t\gg N/b_{1}$). For $N=50$ the onset of $J_{50}$ occurs near $t\approx j_{50,1}/b_{1}\approx 1270$ (the first zero of $J_{50}$ is $j_{50,1}\approx 57.1$), so the domain was extended to $0,\,10^{4}]$; the value is stable to three significant figures for $T\geq 5000$, as confirmed by ARB-certified runs at $T=5000$ ($I=1.4025\times 10^{-14}$) and $T=10^{4}$ ($I=1.4076\times 10^{-14}$).
@@ 6.5. Unconditional finiteness at finite $M$
Remark 6.10 (Subcritical term versus full bound).
Lemma 6.6 bounds $I_{\mathbf{n}}$ by three terms: a subcritical term $C_{1}(e/4)^{|n_{*}|}$ (exponential in $|n_{*}|$) and transition and supercritical terms $C_{2}(T^{*})^{1-(M-|A|)/2}$ and $C_{3}(T^{*})^{1-M/2}$, both of which are polynomial in $|n_{*}|$ since $T^{*}=|n_{*}|/b_{*}$. The proof above uses only the subcritical term. This is justified because for any fixed $M$ and $d_{M}$, the exponential $(e/4)^{N}$ eventually dominates any polynomial in $N$: there exists $N_{0}=N_{0}(M,\{b_{k}\})$ such that $C_{2}N^{1-(M-|A|)/2}\leq(e/4)^{N}$ for all $N>N_{0}$. The finite head $N\leq N_{0}$ contributes a bounded sum regardless of $d_{M}$. Thus $\eta=e/4$ is the eventual geometric rate governing the tail of the resonance sum; it does not claim to be a uniform pointwise bound $I_{\mathbf{n}}\leq C_{0}(e/4)^{N}$ for all $N\geq 1$.
@@
The polynomial decay in $t$ needed for integrability comes from the inactive $J_{0}$ factors: without them the resonance integral decays only as $t^{-|A|/2}$, which cannot overcome even polynomial lattice counting for small $|A|$. The exponential suppression in resonance order $|n_{*}|$ comes instead from the subcritical bound on the active Bessel factor $J_{n_{*}}(b_{*}t)$: the power-series and Stirling estimates give $|J_{n_{*}}(b_{*}t)|\lesssim(e/4)^{|n_{*}|}$ in the subcritical region $b_{*}t<|n_{*}|/2$ (Lemma 6.6). The following table illustrates how the inactive $J_{0}$ factors collectively reduce the magnitude of the resonance integral as $M$ increases:
@@
33 $3.660$ $1.000$
$10$ $1.791$ $0.489$
2020 $1.717$ $0.469$
$30$ $1.703$ $0.465$
Each additional inactive $J_{0}$ factor reduces the resonance integral, converging to a limit as $M\to\infty$. The values of $I_{\mathbf{n}}$ were computed by ARB-rigorous strip decomposition (acb.integral, 256-bit precision, $T=2000$): the domain is partitioned at the certified zeros of each $J_{1}$ and $J_{0}$ factor so that acb.integral is applied only on gap intervals where the integrand is analytic. Active indices $(n_{1},n_{2},n_{3})=(1,1,-1)$; the remaining M3M-3 indices are inactive. All table values are integrals over $[0,2000]$; for $M=3$ (no inactive $J_{0}$ factors) the tail from $2000,\infty)$ is approximately $0.44$ (Watson bound $\leq 2.01$; see ancillary script), so the $M=3$ entry and the ratios normalised by it reflect the truncated integral. For $M\geq 10$ the tail is below $3\times 10^{-5}$ and the table entries are effectively equal to $\int_{0}^{\infty}$.
@@ 7.1. Monotonicity of the main term
The main-term integral is monotone decreasing in $M$.
@@
The main-term sequence $\{f_{\mathrm{main}}^{(M)}\}$ is therefore a decreasing sequence bounded below by zero, and converges to a finite limit. The values in the following table were computed by strip decomposition (§A.6) of $\frac{1}{\pi}\int_{0}^{T}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$ with $T=2000$ using ARB’s acb.integral at 256-bit precision. These are the unsigned main-term integrals and serve as upper bounds on $\|f_{S_{L}^{(M)}}\|_{\infty}$; the signed integrals that equal the actual density $f^{(M)}(0)$ under verified resonance absence are smaller (see Remark 7.2).
@@
33 $12.37$ $7.838$
$10$ $10.585$ $8.292$
2020 $10.517$ $8.313$
This is consistent: each additional $J_{0}(b_{k}t)$ factor reduces the absolute value of the integrand but also reduces the negative lobes, pushing the signed integral upward. The finiteness proof uses only the unsigned bound. The monotone increase of the signed integral confirms that $f^{(M)}(0)$ converges to $f_{S_{L}}(0)=8.3129$ from below. Both columns are the integrals over $[0,T]$ with $T=2000$ at $256$-bit ARB precision; the unsigned column uses strip decomposition (§A.6), the signed column uses direct acb.integral (the integrand $\prod J_{0}$ is analytic). Certified upper bounds on the tail $(1/\pi)\int_{T}^{\infty}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$ via the envelope bound $|J_{0}(x)|\leq\sqrt{2/(\pi x)}$ are reported in the ancillary script; they range from $O(1)$ for $M=3$ to below $10^{-8}$ for $M=20$.
@@ 7.2. The self-referential telescoping bound
In the supercritical region, the uniform asymptotic gives
$|J_{n_{M+1}}(b_{M+1}t)|\leq\sqrt{\frac{2}{\pi\sqrt{(b_{M+1}t)^{2}-n_{M+1}^{2}}}}\leq 2\sqrt{\frac{2}{\pi b_{M+1}t}}$
for $b_{M+1}t\geq 2|n_{M+1}|$. For the remaining product, we use the uniform envelope bound $|J_{n}(x)|\leq\min(1,\sqrt{2/(\pi|x|)})$, which holds for all orders $n\geq 0$ and all $x>0$. This gives
$\prod_{k=1}^{M}|J_{n_{k}}(b_{k}t)|\leq\prod_{k=1}^{M}\min\!\left(1,\sqrt{\frac{2}{\pi b_{k}t}}\right)=:g_{M}(t).$
@@
2020 $1.18\times 10^{-1}$ $9.25\times 10^{-4}$ $128$
$30$ $6.26\times 10^{-2}$ $4.93\times 10^{-4}$ $127$
The ratio $I_{\mathbf{n}}/b_{M}$ is bounded and stabilizing, confirming $C_{M}\leq C$ uniformly. For $n_{M}=2$, the ratio $I_{\mathbf{n}}/b_{M}^{2}$ similarly stabilizes near $1550$. Each entry $I_{\mathbf{n}}$ in the table corresponds to a resonance vector with $n_{M}=1$ and all other components zero; the integral was evaluated by strip decomposition (§A.6; acb.integral on gap intervals, 256-bit precision, $T=2000$).
@@
Since $\prod|J_{0}|$ is integrable on $0,\infty)$ for $M\geq 7$ (the product decays as $t^{-M/2}$, which is integrable when $M\geq 3$; the constant $\frac{1}{4\pi}\int_{0}^{\infty}t^{2}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$ is finite for $M\geq 7$ and decreasing in $M$), the true contribution is $O(b_{M+1}^{2})$ with an effective constant given by $\frac{1}{4\pi}\int_{0}^{\infty}t^{2}\prod_{k=1}^{M}|J_{0}(b_{k}t)|\,dt$.
@@
$10$ $6.699\times 10^{-3}$ $2.468\times 10^{-3}$ $6.092\times 10^{-6}$ $1100$
$15$ $1.827\times 10^{-3}$ $1.375\times 10^{-3}$ $1.892\times 10^{-6}$ $966$
2020 $7.910\times 10^{-4}$ $9.251\times 10^{-4}$ $8.559\times 10^{-7}$ $924$
The ratio $|f^{(M)}(0)-f^{(M-1)}(0)|/b_{M}^{2}$ is bounded and slowly varying, confirming $\alpha=2$. The marginal changes were computed as differences of successive $f^{(M)}(0)$ values obtained by ARB quadrature (acb.integral, 256-bit precision, $T=2000$); the weights $b_{M}=2/(1/4+{\gamma_{M}^{\prime}}^{2})$ use the verified zero ordinates from the appendix.
@@ 7.3. The major-arc/minor-arc decomposition
The values $\epsilon_{M}=\sum_{k>M}b_{k}$ below are computed from the first 200 certified zeros of $L(s,\chi_{5})$ (Appendix C) plus the analytic tail bound $0.0071$ via Abel summation against the zero-counting formula; see Appendix A for the full methodology.
@@
Under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (computationally consistent with $d_{M}=0$ for $M\leq 20$; the hypothesis asserts this extends to the full infinite lattice), the Cesàro distribution of $S_{L}(\delta)=\sum_{k=1}^{\infty}b_{k}\cos(\gamma_{k}^{\prime}\delta)$ has a density at the origin satisfying $f_{S_{L}}(0)<\infty$. The same holds for $f_{S_{\zeta}}(0)$.
@@
Can the passage from verified resonance absence at finite $M$ to the limiting density $f_{S_{L}}(0)<\infty$ be achieved by purely approximation-theoretic means, bypassing the hypothesis $d_{\infty}<\infty$? Since $S_{L}^{(M)}\to S_{L}$ in $B^{2}$, the value distributions converge weakly (by Cauchy–Schwarz applied to characteristic functions and Lévy continuity). However, weak convergence does not imply pointwise convergence of densities. The Fourier inversion formula $f^{(M)}(0)=(2\pi)^{-1}\int\prod_{k=1}^{M}J_{0}(b_{k}t)\,dt$ requires the product factorization of the characteristic function, which in turn requires $\mathbb{Q}$-linear independence of the frequencies $\gamma_{1}^{\prime},\ldots,\gamma_{M}^{\prime}$. The dominated convergence argument that yields $f^{(M)}(0)\to f^{(\infty)}(0)$ relies on the monotone bound $|\varphi^{(M+1)}(t)|\leq|\varphi^{(M)}(t)|$, which holds only under the product structure. Without independence, the characteristic function need not factor, the monotone domination fails, and no alternative $L^{1}$ dominator is available. The obstruction is therefore structural, not computational: every Fourier-analytic path to $f_{S_{L}}(0)<\infty$ passes through the independence of zero ordinates.